# Generalized Bayesian deep reinforcement learning

### v1

@@ 3.2 Prequential scoring rule posterior
    $\operatorname{\mathcal{PS}}_{T}^{a}:\Theta\rightarrow\mathbb{R}$ has continuous and uniformly bounded third derivatives on $E$, and if $T\rightarrow\infty$ then $\frac{\operatorname{\mathcal{PS}}_{T}^{a}}{|T_{a}|}\rightarrow\xi^{a}$ pointwise for some $\xi^{a}:\Theta\rightarrow\mathbb{R}$ with positive definite second derivative at $\theta^{*}$, denoted as $\operatorname{\Gamma}^{a}_{\star}$, $\forall a\in\mathcal{A}$.
@@
    For some compact set $K\subseteq E$ with $\theta^{*}$ in its interior, $\xi^{a}(\theta)>\xi^{a}\left(\theta^{*}\right)$ for all $\theta\in K\backslash\left\{\theta^{*}\right\}$ and $\liminf_{T}\inf_{\theta\in\Theta\backslash K}\operatorname{\mathcal{PS}}_{T}^{a}(\theta)>\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta^{*}\right)$, $\forall a\in\mathcal{A}$.
@@
    $\operatorname{\mathcal{PS}}_{T}^{a}$ is convex and $\xi^{a\prime}\left(\theta^{*}\right)=0$, $\forall a\in\mathcal{A}$.
@@
If A1 and (A2 or A3) holds for all actions $a\in\mathcal{A}$, then the total prequential Scoring Rule loss, $\operatorname{\mathcal{PS}}_{T}(\theta)=\sum_{a\in\mathcal{A}}\operatorname{\mathcal{PS}}_{T}^{a}(\theta)$ also satisfies these conditions and there exists a sequence $\theta_{T}$ that converges to $\theta^{*}$ almost surely as $T\rightarrow\infty$.
We do not provide a proof as a similar result for generalized posteriors is established in Theorem 5 of Miller (2021). Note that, if the above conditions hold for each action-specific prequential SR $\operatorname{\mathcal{PS}}^{a}$, then the conditions are also satisfied by the summation $\operatorname{\mathcal{PS}}$. Consequently, Theorem 5 in Miller (2021) can be extended to prove the existence of a sequence of consistent estimators of $\theta^{*}$ as stated in Lemma 1. This result is analogous to the consistency of maximum likelihood estimators (MLE) and by setting $S$ as the log score, the original result can be recovered.
Next we prove the BvM type result for prequential scoring rules for MDPs under three additional assumptions. Note that assumption A4 ensures a polynomial approximation of the prequential SRs through a Taylor series expansion. Whereas assumption A6 ensures that all the actions are well explored in the training dataset. Under these assumptions, we formally state the theorem bellow.
@@
    $\operatorname{\mathcal{PS}}_{T}^{a}$ can be represented as
    $\frac{1}{|T_{a}|}\operatorname{\mathcal{PS}}_{T}^{a}(\theta)=\frac{1}{|T_{a}|}\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta_{T}\right)+\frac{1}{2}\left(\theta-\theta_{T}\right)^{\mathrm{T}}\operatorname{\Gamma}_{T}^{a}\left(\theta-\theta_{T}\right)+r_{T}^{a}\left(\theta-\theta_{T}\right),$
    where $\theta_{T}\in\mathbb{R}^{d}$ is such that $\theta_{T}\rightarrow\theta^{*}$ and $\operatorname{\Gamma}_{T}^{a}\in\mathbb{R}^{d\times d}$ is symmetric such that $\operatorname{\Gamma}_{T}^{a}\rightarrow\operatorname{\Gamma}_{*}^{a}$ as $|T_{a}|\rightarrow\infty$, $\forall a\in\mathcal{A}$. Further, there exist $\varepsilon_{0},c_{0}>0$ such that for all $T$ sufficiently large, for all $x\in B_{\varepsilon_{0}}(0)$, $\forall a\in\mathcal{A}$, $r_{T}^{a}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfies $\left|r_{T}^{a}(x)\right|\leq c_{0}|x|^{3}$.
@@
    For any $\varepsilon>0,\liminf_{n}\inf_{\theta\in B_{\varepsilon}\left(\theta_{T}\right)^{c}}\frac{1}{|T_{a}|}\left(\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta_{T}\right)\right)>0$ $\forall a\in\mathcal{A}$.
@@
    $\lim_{\begin{subarray}{c}|T_{a}|\rightarrow\infty\\ T\rightarrow\infty\end{subarray}}\frac{|T_{a}|}{T}=w_{a}\in(0,1)$, $\forall a\in\mathcal{A}$.
@@
that is, $\pi_{\operatorname{\mathcal{PS}}}(\theta|H_{T})$ concentrates at $\theta^{*}$,
$z_{T}\sim\frac{\exp\left(-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)\right)\pi\left(\theta^{*}\right)}{\left|\operatorname{det}\operatorname{\Gamma}_{*}\right|^{1/2}}\left(\frac{2\pi}{T}\right)^{D/2}$ (12)
as $T\rightarrow\infty$ (Laplace approximation), with $\operatorname{\Gamma}_{*}=\sum_{a}w_{a}\operatorname{\Gamma}_{*}^{a}$; and letting $q_{T}$ be the density of $\sqrt{T}\left(\theta-\theta_{T}\right)$ when $\theta\sim\pi_{\operatorname{\mathcal{PS}}}\left(\theta\mid H_{T}\right)$,
$\int_{\mathbb{R}^{d}}\left|q_{T}(x)-\mathcal{N}\left(x\mid 0,\operatorname{\Gamma}_{*}^{-1}\right)\right|dx{\longrightarrow}0$ (13)
as $T\rightarrow\infty$, that is, $q_{T}$ converges almost surely to $\mathcal{N}\left(0,\operatorname{\Gamma}_{*}^{-1}\right)$ in total variation.
The above theorem implies that, as the number of observations, $T$, increases, the scoring rule posterior derived from these observations increasingly resembles a Gaussian distribution. This posterior distribution becomes progressively more concentrated around the expected scoring rule minimizer. We provide the full proof of Theorem 2 in Appendix A.1.
3.3 Sequential Monte Carlo with gradient based kernel
@@
Sequential Monte Carlo (SMC) refers to a class of algorithms which aim to represent a probability distribution through a set of weighted particles. Let $\{\pi_{k}\}$, for $k=1,2,\ldots$, to be the sequence of scoring rule posteriors using data from all the episodes up to the $k$-th episode. For each $k$, we use SMC to sample from the corresponding $\pi_{k}$, initializing the SMC with the posterior samples generated from the scoring rule posterior of the $(k-1)$-th episode. Hence the prior distribution used at the $k$-th episode to define the scoring rule posterior is $\pi_{k-1}$. To ensure a smooth transition from the prior ($\pi_{k-1}$) to the target distribution ($\pi_{k}$) at each episode, we introduce a series of intermediate target distributions. These distributions follow a geometric path, defined as ${\pi}^{(l)}_{k}(\theta)\propto\pi^{\alpha_{k,l}}_{k}(\theta)\pi_{k-1}^{1-\alpha_{k,l}}(\theta)$ with $10\leq\alpha_{k,1}<\ldots\alpha_{k,L}=$ 1 as proposed by Gelman and Meng (1998). The sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode can be determined adaptively based on effective sample size (ESS)(Beskos et al., 2015) or conditioned effective sample size (CESS)(Zhou et al., 2016). Note that the prior distribution $\pi_{0}$ for the first episode, preferably an informative one, needs to be chosen.
The core of SMC relies on importance sampling across the particle trajectories. However, due to the high computational cost associated with sequential importance sampling, SMC introduces a series of backward kernels, used alongside forward kernels. Starting from an initial set of weighted particles drawn from a proposal distribution, each SMC iteration propagates these particles toward the target distribution via a forward kernel. The resulting trajectories are reweighted. The particles are then resampled according to these updated (normalized) weights. Note that we used an MCMC kernel with an invariant distribution matching the target posterior as the forward kernel and the corresponding time-reversal kernel as the backward kernel as outlined in Section 3.3.2.3 in Del Moral et al. (2006).
Before providing the details of the gradient based kernel, we first argue how we can easily derive an unbiased estimate of the prequential scoring rule. As discussed in Section 3.1, the prequential SR $S({P}_{\theta_{st}}(.|s_{t},a_{t}),s_{t+1})$ can also be expressed as an expectation over samples from $P_{\theta_{st}}$, conditioned on $(s_{t},a_{t})$ for all $t=1,2,\ldots T$. Whenever $S$ is differentiable with respect to $\theta$, an unbiased estimator of the gradient of the total prequential SR $\operatorname{\mathcal{PS}}_{T}(\theta)$ can be obtained using random samples from an auxiliary distribution such as a Gaussian or uniform distribution that is independent of $\theta$ using equation (5). Next, we describe the adSGRLD kernel used as the forward kernel for SMC.
@@
Suppose we wish to sample from a distribution with pdf $p(\theta)=(1/Z)\exp(-U(\theta))$, $\theta\in\mathbb{R}^{d}$; where $Z$ is a normalizing constant and $U(\theta)$, the negative log-likelihood equivalent, is known as the potential energy. Standard SGLD is based on the Overdamped Langevin Diffusion, represented by the following stochastic differential equation which has a stationary distribution $p\left(\theta\right)$:
@@
where $B_{u}$ is a Brownian motion. To sample from the target distribution using the above SDE, we often use numerical approximation schemes like the Euler-Maruyama discretisation, which leads to the following update rule:
$\theta_{u+1}\leftarrow\theta_{u}-\frac{\epsilon}{2}\widehat{\nabla_{\theta}}U(\theta_{u})+\sqrt{\epsilon_{u}}W,$
@@
where $W$ is a d-dimensional standard normal random vector; $\widehat{\nabla_{\theta}}U(\theta_{u})$ is an unbiased estimator of the gradient of $U(\theta_{u})$ and $\{\epsilon_{u}\}$ is a sequence of discretization step sizes satisfying the conditions, $\sum_{u=1}^{\infty}\epsilon_{u}=\infty$ and $\sum_{u=1}^{\infty}\epsilon_{u}^{2}<\infty$.
@@
Although Welling and Teh (2011) have shown that SGLD yields samples from the target posterior when using a sequence $\epsilon_{u}$ converging to 00, in practice, $\epsilon_{u}$ seldom converges to 00, leading to bias due to Euler-Maruyama discretization. Hence, to ensure random sampling with minimal bias even when using a noisy estimate of the gradient, adaptive Langevin dynamics has been proposed in Jones and Leimkuhler (2011) which was later adapted for Bayesian inference (Ding et al., 2014) and likelihood-free inference (Pacchiardi et al., 2024b). The algorithm, referred to as adaptive stochastic gradient Langevin dynamics (adSGLD), runs on an augmented space $(\theta,p,\eta)$, where $\theta$ represents the parameter of interest, $p\in\mathbb{R}^{d}$ represents the momentum and $\eta$ represents an adaptive thermostat controlling the mean kinetic energy $\frac{1}{d}\mathbb{E}(p^{T}p)$.
@@
Choosing an appropriate sequence of step sizes is crucial for effectively exploring the parameter space, especially in high dimensions. For example if different components of $\theta$ have values on different scales or the components are highly correlated, a poor choice of step sizes can result in slow mixing, negatively impacting the performance of the SMC sampler. To address this issue, several preconditioning schemes (Girolami and Calderhead, 2011) have been proposed in the literature. For instance, the Riemann manifold Metropolis-adjusted Langevin algorithm uses a positive definite matrix $G(\theta)$ to adaptively precondition the gradient. Then the SDEs are given by,
$d\bm{\theta}$ $=$ $G(\theta)\bm{p}du,$
$d\bm{p}$ $=$ $\left(-G(\bm{\theta})\nabla_{\bm{\theta}}U(\bm{\theta})-\bm{\eta}\bm{p}+\nabla_{\bm{\theta}}G(\bm{\theta})+G(\bm{\theta})(\bm{\eta}-G(\bm{\theta}))\nabla_{\bm{\theta}}G(\bm{\theta})\right)du+\sqrt{2}G(\bm{\theta})^{\frac{1}{2}}dB_{u},$
$d\bm{\eta}$ $=$ $\left(\frac{1}{d}\bm{p}^{T}\bm{p}-1\right)du$
@@
where $G(\theta)$ is our preconditioning matrix, and $\nabla_{\bm{\theta}}G(\bm{\theta})$ is a vector with $i$-th element being $\sum_{j}\nabla_{\bm{\theta}_{j}}G_{ij}(\bm{\theta})$. $G(\bm{\theta})$ encodes geometric information of the potential energy $U(\bm{\theta})$, called Riemannian metric (Girolami and Calderhead, 2011), which are commonly defined by the Fisher information matrix.
@@
If we assume $G(\bm{\theta})$ does not depend on $\bm{\theta}$, we notice that the third and fourth terms in the momentum update term vanishes and we end up with a SDE of the form
$d\bm{\theta}$ $=$ $G(\theta)\bm{p}du,$ (14)
$d\bm{p}$ $=$ $\left(-G(\bm{\theta})\nabla_{\bm{\theta}}U(\bm{\theta})-\bm{\eta}\bm{p}\right)du+\sqrt{2}G(\bm{\theta})^{\frac{1}{2}}dB_{u},$
$d\bm{\eta}$ $=$ $\left(\frac{1}{d}\bm{p}^{T}\bm{p}-1\right)du$
Using an Euler scheme with step size $\epsilon$, our algorithm becomes
@@
Initialize $\bm{\theta}_{(0)}\in\mathbb{R}^{n},\mathbf{p}_{(0)}\sim\mathcal{N}(0,G(\bm{\theta}_{(0)}\epsilon\mathbf{I})$, and $\bm{\eta}_{(0)}=a$

### v2

@@ 3.2 Prequential scoring rule posterior
    $S$ is a strictly proper scoring rule and, the time-averaged generalized entropy of the true model $\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{0},\mathbb{M}_{0})$ has a finite limit as $T\to\infty$.
Lemma 1
Under Assumption A1, there exists a function $\operatorname{\mathcal{PS}}^{*}(\theta)$ such that as $T\to\infty$, $\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)\to\operatorname{\mathcal{PS}}^{*}(\theta)$ uniformly with probability one under $\mathbb{M}_{0}$.
An outline of the proof of the Lemma 1 is provided in Appendix A.1. Next, let us denote the minimizer of the (normalized) empirical prequential score as
$\hat{\theta}_{T}=\arg\min_{\theta\in\Theta}\frac{1}{T}\operatorname{\mathcal{PS}}_{T}(\theta).$
A key step in establishing the consistency of this estimator is a uniform law of large numbers (ULLN) for the empirical prequential scores. However, classical ULLNs assume i.i.d. observations, which do not apply here as the data are generated by a Markov process. We therefore impose mixing conditions and exploit an action-based decomposition of the prequential score to obtain a ULLN.
Now suppose the action space $\mathcal{A}$ is finite. Then, we can decompose the prequential score as,
$\mathcal{PS}_{T}(\theta)=\sum_{a\in\mathcal{A}}\mathcal{PS}_{T}^{a}(\theta),$
where $\mathcal{PS}_{T}^{a}(\theta)=\operatorname{\mathcal{PS}}_{T}(\mathbb{M}_{\theta},h_{T}^{a})=\sum_{t=1}^{T}S({P}_{\theta}(.|s_{t},a_{t}),s_{t+1})\mathbb{I}_{\{a_{t}=a\}}$ records the scoring rule only for the time-steps at which action $a$ was taken upto time $T$. In addition, $T_{a}$ denotes the set of such time points, i.e. $T_{a}=\{t\leq T:a_{t}=a\}$ for each $a\in\mathcal{A}$. So, on fixing the current action, the simulator model $P_{\theta}$ merely predicts $s_{t+1}$ conditioned on the current state $s_{t}$. We use this decomposition along with the following assumptions on the true data generating process $\mathbb{M}_{0}$ to establish a ULLN.
@@
    (Asymptotic stationarity) For all actions $a$ in the action space, let, $G_{t}^{a}$ be the marginal distribution of $(S_{t},S_{t+1})$ when $A_{t}=a$ for $t\geq 1$; then, $\frac{1}{|T_{a}|}G_{t}^{a}$ converges weakly to some probability measure $G^{a}$ on $\mathcal{S}^{2}$ as $T\to\infty$.
@@
    For all actions $a$ in the action space, both of the conditions below are satisfied:
    1. (a)
      (Mixing) Suppose $(S_{t})_{t}^{a}$ is the sequence of the states at which action $a$ is taken. Either one of the following conditions holds:
      1. i.
        $(S_{t})_{t}^{a}$ is $\alpha$-mixing with mixing coefficient of size $r/(2r-1)$, with $r\geq 1$, or
      2. ii.
        $(S_{t})_{t}^{a}$ is $\gamma$-mixing with mixing coefficient of size $r/(r-1)$, with $r>1$.
    2. (b)
      (Moment boundedness) Define $J^{a}(s_{t},s_{t+1})=\sup_{\theta\in\Theta}|S(P_{\theta}(\cdot|s_{t},a),s_{t+1})|$, then
      $\sup_{t\geq 1}\mathbb{E}[J^{a}(s_{t},s_{t+1})^{r+\delta}]<\infty,$
      for some $\delta>0$, for the value of $r$ corresponding to the condition above which is satisfied.
@@
(Uniform law of large numbers) Under assumptions A2 and A3, a ULLN holds, which implies with probability $1$ under $\mathbb{M}_{0}$,
$\sup_{\theta\in\Theta}|\operatorname{\mathcal{PS}}_{T}(\theta)-\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)|\to 0\text{ as }T\to\infty.$
Proof of Lemma 2 is provided in Appendix A.2. We use a result from Pötscher and Prucha (1989) as adapted by Pacchiardi et al. (2024a) to establish the ULLN for prequential scores under the true data-generating process, assuming $\mathbb{M}_{0}$ satisfies the asymptotic stationarity and mixing conditions stated in Assumptions A2 and A3. Intuitively, asymptotic stationarity implies that the joint distribution of any two consecutive states, conditioned on an action, converges to a stationary distribution. The mixing conditions ensure that, under a fixed action, the dependence between states decays rapidly as their time separation increases. These properties automatically hold for the Markovian model $\mathbb{M}_{0}$, where, given an action, the next state depends only on the current state. However, we prove the result for a broader class of models which satisfy the conditions A2 and A3.
Corollary 3
Under assumptions A1-A3, it follows from Lemma 1 and Lemma 2 that,
$\sup_{\theta\in\Theta}|\frac{1}{T}\operatorname{\mathcal{PS}}_{T}(\theta)-{\operatorname{\mathcal{PS}}}^{*}(\theta)|\to 0\text{ as }T\to\infty,$
with probability $1$ under $\mathbb{M}_{0}$.
The above corollary is a direct consequence of Lemma 1 and Lemma 2 and a short proof can be found in Appendix A.3. Next, we use this result to show the asymptotic consistency of the estimators $\{\hat{\theta}_{T}\}_{T=1}^{\infty}$ under the following assumptions.
@@
    The parameter space $\Theta\subseteq\mathbb{R}^{p}$ is compact.
@@
    $\theta^{*}$ is a unique minimizer of $\operatorname{\mathcal{PS}}^{*}(\theta)$, and there exists a metric $d$ on $\Theta$ such that, for all $\epsilon>0,$
    $\min_{\theta:d(\theta,\theta^{*})\geq\epsilon}{\operatorname{\mathcal{PS}}}^{*}(\theta)-{\operatorname{\mathcal{PS}}}^{*}(\theta^{*})>0.$
Lemma 4
(Asymptotic consistency) Under assumptions A1-A5, as $T\to\infty$, we have with probability $1$ under $\mathbb{M}_{0}$,
$d(\hat{\theta}_{T},\theta^{*})\to 0.$
We derive the above lemma using a result from Skouras (1998) and a detailed proof can be found in Appendix A.4. Here the parameter space needs to be compact and the empirical prequential score must be smooth enough in the neighbourhood of $\theta^{*}$ as in condition A5. Then, this result is analogous to the consistency of maximum likelihood estimators (MLE). Specifically, when the scoring rule $S$ is the log score, the estimator $\hat{\theta}_{T}$ is the MLE. The result establishes that $\hat{\theta}_{T}$ is consistent for $\theta^{*}$, the true parameter under a well-specified model, i.e., when $\mathbb{M}_{0}=\mathbb{M}_{\theta^{*}}$. In the misspecified case, where $\mathbb{M}_{0}\notin\mathcal{M}$, the parameter $\theta^{*}$ corresponds to the model within $\mathcal{M}$ that minimizes the expected prequential loss—i.e., the model with optimal one-step-ahead predictive performance. Using these consistent estimators of $\theta^{*}$, we next state the BvM theorem for the generalized prequential posterior under the following assumptions.
@@
    $\pi:\mathbb{R}^{p}\rightarrow\mathbb{R}$ is a probability distribution with respect to the Lebesgue measure such that $\pi$ is continuous at $\theta^{*}$ and $\pi(\theta^{*})>0$.
  2. A7
    $E\subseteq\mathbb{R}^{p}$ is open and convex and let $\theta^{*},\hat{\theta}_{T}\in E$ for all $T$ sufficiently large.
  3. A8
    $\operatorname{\mathcal{PS}}_{T}^{\prime\prime}(\theta^{*})\to H^{*}$ as $T\to\infty$ for some positive definite $H^{*}$.
  4. A9
    $\operatorname{\mathcal{PS}}_{T}(\theta)$ have continuous third derivatives in $E$ and the third derivatives $\operatorname{\mathcal{PS}}^{{}^{\prime\prime\prime}}_{T}(\theta)$ are uniformly bounded in $E$.
  5. A10
    For any $\epsilon>0$, $\lim\inf_{T}\inf_{\theta:d(\theta,\hat{\theta}_{T})\geq\epsilon}(\operatorname{\mathcal{PS}}_{T}(\theta)-\operatorname{\mathcal{PS}}_{T}(\hat{\theta}_{T}))>0.$
@@ 5. A10
that is, $\pi_{\operatorname{\mathcal{PS}}}(\theta|h_{T})$ concentrates at $\theta^{*}$;
$z_{T}\sim\frac{\exp\left(-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)\right)\pi\left(\theta^{*}\right)}{\left|\operatorname{det}H^{*}\right|^{1/2}}\left(\frac{2\pi}{T}\right)^{p/2}$ (11)
as $T\rightarrow\infty$ (Laplace approximation); and letting $q_{T}$ be the density of $\sqrt{T}\left(\theta-\hat{\theta}_{T}\right)$ when $\theta\sim\pi_{\operatorname{\mathcal{PS}}}\left(\theta\mid h_{T}\right)$,
$\int_{\mathbb{R}^{D}}\left|q_{T}(\theta)-\mathcal{N}\left(\theta\mid 0,H^{*-1}\right)\right|d\theta{\longrightarrow}0$ (12)
as $T\rightarrow\infty$, that is, $q_{T}$ converges almost surely to $\mathcal{N}\left(0,H^{*-1}\right)$ in total variation.
We prove the above BvM theorem using two results from Miller (2021), and we provide a sketch of the proof in Appendix A.5. For this result to hold, the prior distribution must be continuous and positive at $\theta^{*}$. Both $\theta^{*}$ and the sequence of its consistent estimators $\hat{\theta}_{T}$ must lie in an open and convex subset of $\Theta$. Further, the assumptions A8 and A9 ensure a Taylor series expansion of the empirical prequential score $\operatorname{\mathcal{PS}}_{T}(\theta)$. In addition, we assume that $\operatorname{\mathcal{PS}}_{T}(\theta)$ is smooth in the neighbourhood of $\hat{\theta}_{T}$ as in condition A10. Then, the theorem states that as more data is observed, the generalized prequential scoring rule posterior concentrates around the $\theta^{*}$. It also begins to resemble a Gaussian distribution. The covariance matrix of this Gaussian is given by the inverse of $H^{*}$ which is a limit of $\operatorname{\mathcal{PS}}^{\prime\prime}_{T}(\theta)$. Hence, $H^{*}$ plays the role of the Fisher information matrix in the likelihood-based framework.
3.3 Sequential Monte Carlo with gradient-based kernel
@@
Sequential Monte Carlo (SMC) refers to a class of algorithms that aim to represent a probability distribution through a set of weighted particles. Let $\{\pi_{k}\}$, for $k=1,2,\ldots$, to be the sequence of scoring rule posteriors using data from all the episodes up to the $k$-th episode. For each $k$, we use SMC to sample from the corresponding $\pi_{k}$, initializing the SMC with the posterior samples generated from the scoring rule posterior of the $(k-1)$-th episode. Hence the prior distribution used at the $k$-th episode to define the scoring rule posterior is $\pi_{k-1}$. To ensure a smooth transition from the prior ($\pi_{k-1}$) to the target distribution ($\pi_{k}$) at each episode, we introduce a series of intermediate target distributions. These distributions follow a geometric path, defined as ${\pi}^{(l)}_{k}(\theta)\propto\pi^{\alpha_{k,l}}_{k}(\theta)\pi_{k-1}^{1-\alpha_{k,l}}(\theta)$ with $10\leq\alpha_{k,1}<\ldots\alpha_{k,L}=$ 1 as proposed by Gelman and Meng (1998). The sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode can be determined adaptively based on effective sample size (ESS) (Beskos et al., 2015) or conditional effective sample size (CESS) (Zhou et al., 2016).
An SMC sampler uses sequential importance sampling with a resampling approach on the sequence of targets. Starting from an initial set of weighted particles drawn from a proposal distribution, each SMC iteration propagates these particles toward the target distribution via a forward kernel. The resulting trajectories are reweighted. The particles are then resampled according to these updated (normalized) weights. Note that we used an MCMC kernel with an invariant distribution matching the target posterior as the forward kernel. SMC samplers also require the specification of backward kernels. We use the corresponding time-reversal kernel as the backward kernel as outlined in Section 3.3.2.3 in Del Moral et al. (2006).
Before providing the details of the gradient-based kernel, we first argue how we can easily derive an unbiased estimate of the prequential scoring rule. As discussed in Section 3.1, the prequential SR $S({P}_{\theta_{st}}(.|s_{t},a_{t}),s_{t+1})$ can also be expressed as an expectation over samples from $P_{\theta_{st}}$, conditioned on $(s_{t},a_{t})$ for all $t=1,2,\ldots T$. Whenever $S$ is differentiable with respect to $\theta$, an unbiased estimator of the gradient of the total prequential SR $\operatorname{\mathcal{PS}}_{T}(\theta)$ can be obtained using random samples from an auxiliary distribution such as a Gaussian or uniform distribution that is independent of $\theta$ using equation (6). Next, we describe the adjusted stochastic gradient Riemannian Langevin dynamic (adSGRLD) kernel used as the forward kernel for SMC.
@@
Suppose we wish to sample from a distribution with pdf $p(\theta)=(1/Z)\exp(-U(\theta))$, $\theta\in\mathbb{R}^{p}$; where $Z$ is a normalizing constant and $U(\theta)$, the negative log-likelihood equivalent, is known as the potential energy. Standard SGLD is based on the Overdamped Langevin Diffusion, represented by the following stochastic differential equation which has a stationary distribution $p\left(\theta\right)$:
@@
where $B_{u}$ is a Brownian motion. To sample from the target distribution using the above SDE, we often use numerical approximation schemes like the Euler-Maruyama discretization, which leads to the following update rule:
$\theta_{u+1}\leftarrow\theta_{u}-\frac{\epsilon}{2}\widehat{\nabla_{\theta}}U(\theta_{u})+\sqrt{\epsilon_{u}}W,$ (13)
@@
where $W$ is a d-dimensional standard normal random vector; $\widehat{\nabla_{\theta}}U(\theta_{u})$ (often used in practice) is an unbiased estimator of the gradient of $U(\theta_{u})$ and $\{\epsilon_{u}\}$ is a sequence of discretization step sizes satisfying the conditions, $\sum_{u=1}^{\infty}\epsilon_{u}=\infty$ and $\sum_{u=1}^{\infty}\epsilon_{u}^{2}<\infty$.
@@
Although Welling and Teh (2011) have shown that SGLD yields samples from the target posterior when using a sequence $\epsilon_{u}$ converging to 00, in practice, $\epsilon_{u}$ seldom converges to 00, leading to bias due to Euler-Maruyama discretization. Hence, to ensure random sampling with minimal bias even when using a noisy estimate of the gradient, adaptive Langevin dynamics has been proposed in Jones and Leimkuhler (2011) which was later adapted for Bayesian inference (Ding et al., 2014) and likelihood-free inference (Pacchiardi et al., 2024b). The algorithm, referred to as adaptive stochastic gradient Langevin dynamics (adSGLD), runs on an augmented space $(\theta,\kappa,\eta)$, where $\theta$ represents the parameter of interest, $\kappa\in\mathbb{R}^{p}$ represents the momentum and $\eta$ represents an adaptive thermostat controlling the mean kinetic energy $\frac{1}{p}\mathbb{E}(\kappa^{T}\kappa)$.
@@
Choosing an appropriate sequence of step sizes is crucial for effectively exploring the parameter space, especially in high dimensions. For example, if different components of $\theta$ have values on different scales or the components are highly correlated, a poor choice of step sizes can result in slow mixing, negatively impacting the performance of the SMC sampler. To address this issue, several preconditioning schemes (Girolami and Calderhead, 2011) have been proposed in the literature. For instance, the Riemann manifold Metropolis-adjusted Langevin algorithm uses a positive definite matrix $G(\theta)$ to adaptively precondition the gradient. Then the SDEs are given by,
$d\bm{\theta}$ $=$ $G(\theta)\bm{\kappa}du,$
$d\bm{\kappa}$ $=$ $\left(-G(\bm{\theta})\nabla_{\bm{\theta}}U(\bm{\theta})-\bm{\eta}\bm{\kappa}+\nabla_{\bm{\theta}}G(\bm{\theta})+G(\bm{\theta})(\bm{\eta}-G(\bm{\theta}))\nabla_{\bm{\theta}}G(\bm{\theta})\right)du+\sqrt{2}G(\bm{\theta})^{\frac{1}{2}}dB_{u},$
$d\bm{\eta}$ $=$ $\left(\frac{1}{p}\bm{\kappa}^{T}\bm{\kappa}-1\right)du$
@@
where $G(\theta)$ [^4] We used the preconditioning matrix inspired by Adam optimizers as suggested in Chen et al. (2016). is our preconditioning matrix, and $\nabla_{\bm{\theta}}G(\bm{\theta})$ is a vector with $i$-th element being $\sum_{j}\nabla_{\bm{\theta}_{j}}G_{ij}(\bm{\theta})$. $G(\bm{\theta})$ encodes geometric information of the potential energy $U(\bm{\theta})$, called Riemannian metric (Girolami and Calderhead, 2011), which are commonly defined by the Fisher information matrix.
@@
If we assume $G(\bm{\theta})$ does not depend on $\bm{\theta}$, we notice that the third and fourth terms in the momentum update term vanishes and we end up with an SDE of the form
$d\bm{\theta}$ $=$ $G(\theta)\bm{\kappa}du,$
$d\bm{\kappa}$ $=$ $\left(-G(\bm{\theta})\nabla_{\bm{\theta}}U(\bm{\theta})-\bm{\eta}\bm{\kappa}\right)du+\sqrt{2}G(\bm{\theta})^{\frac{1}{2}}dB_{u},$
$d\bm{\eta}$ $=$ $\left(\frac{1}{p}\bm{\kappa}^{T}\bm{\kappa}-1\right)du$ (14)
Using an Euler scheme with step size $\epsilon$ in the set of Equations 3.3, we obtain the following sampling Algorithm 2 where in place of $\nabla_{{\theta}}U({\theta})$, we use its unbiased estimate $\widehat{\nabla_{\theta}}U(\theta_{u})$ (similar to equation (13)).
@@
Initialize $\bm{\theta}_{(0)}\in\mathbb{R}^{p},\mathbf{\kappa}_{(0)}\sim\mathcal{N}(0,G(\bm{\theta}_{(0)}\epsilon\mathbf{I})$, and $\bm{\eta}_{(0)}=a$